OpenStudy (anonymous):

write the equation x^2 + 4y^2 - 6x - 7=0 in the standard form. state whether the graph is a parabola, circle, ellipse, or hyperbola.

OpenStudy (anonymous):

It's an ellipse

OpenStudy (anonymous):

thanks what about this one? 7x^2 - 28x + 4y^2 - 8y= -2

OpenStudy (anonymous):

\[x^{2} + 4y^{2} - 6x = 7\]\[x^{2} - 6x + 9 + 4y^{2} = 16\]\[\frac{\left( x - 3 \right)^{2}}{16} + \frac{y^{2}}{4} = 1\]the graph is an ellipse